OpenStudy (xixi743):

Find the linearization of L(x) of a function at a. f(x) = e * ln(x^3), a=e^2

OpenStudy (xixi743):

original equation:\[e*\ln(x^3), a = e^2\]

OpenStudy (baru):

\(\Delta y=f'(a) \Delta x\)